OpenStudy (anonymous):

Enter the expression 2cos2(θ)−1, where θ is the lowercase Greek letter theta.

OpenStudy (sidsiddhartha):

its a basic identity \[2 \cos^2 \theta-1=\cos 2 \theta\]